# Cost-aware Portfolios in a Large Universe of Assets

### v1

@@ 2.3.1 First Portfolio Construction
$\min_{\bm{w}_{1}}\bm{w}^{\top}_{1}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}\text{ s.t. }\bm{w}_{1}^{\top}\mathbf{1}=1.$ (2.13)
where $\bm{\alpha}=(\alpha_{1},...,\alpha_{p})^{\top},\alpha_{j}>0$ is the cost parameter, which is similar to $\bm{\beta}$ in Section 2.2. The first-order condition of the constrained optimization (2.13) is
$2\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{\mu}_{1}+\bm{\alpha}\odot\mathbf{g}-h^{*}\mathbf{1}=0,$ (2.14)
where $\mathbf{g}$ is the vector of sub-derivatives of $\left\|\bm{w}_{1}\right\|_{1}$, i.e., $\mathbf{g}:=\partial\left\|\bm{w}_{1}\right\|_{1}/\partial\bm{w}_{1}$, consisting of elements which are 1 or $-1$ in case $w_{j}>0$ or $w_{j}<0$, respectively, or $\in[-1,1]$ in case $w_{j}=0$, and $h^{*}=\frac{2-\gamma\mathbf{1}^{\top}\bm{\Sigma}_{1}^{-1}\bm{\mu}_{1}+\mathbf{1}^{\top}\bm{\Sigma}_{1}^{-1}(\bm{\alpha}\odot\mathbf{g})}{\mathbf{1}^{\top}\bm{\Sigma}_{t}^{-1}\mathbf{1}}$ denotes the Lagrangian multiplier. By Lagrange multiplier method and the definition of $\bm{w}^{*}_{1}$ and $\mathcal{A}_{1}$, solving for $\bm{w}_{1}^{*}$ yields
$\bm{w}_{1}^{*}=\frac{1}{2}\bm{\Sigma}^{-1}_{t}\left(\gamma\bm{\mu}_{1}-\bm{\alpha}\odot\mathbf{g}+\frac{2-\gamma\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{1}\bm{\mu}_{1}+\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{1}(\bm{\alpha}\odot\mathbf{g})}{\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{1}\mathbf{1}}\mathbf{1}\right)\quad\text{and}\quad\bm{w}^{*}_{\mathcal{A}^{c}_{1}}=\mathbf{0}.$
In high-dimensional regimes, we consider the following constrained and regularized quadratic programming to encourage a sparse estimator of $\bm{w}^{*}_{1}$:
$\min_{\bm{w}_{1}}\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}+P_{\lambda}(\bm{w}_{1}),\text{ s.t. }\bm{w}_{1}^{\top}\mathbf{1}=1.$ (2.15)
Similar to Lemma 1 in Section 2.2, Lemma 2 shows the consistency of the initial Lasso weight estimator under proportional transaction cost.
@@
$\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{w}_{1}=1}\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}+\lambda_{\text{Lasso}}\|\bm{w}_{1}\|_{1},$ (2.16)
with $\lambda_{\text{Lasso}}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$, then with probability at least $1-c_{1}p^{-c_{2}}$ for some constants $c_{1},c_{2}$, we have
$\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}-\bm{w}_{1}^{*}\|_{1}=O\left(s_{1}(\alpha+\sqrt{\frac{\log p}{n}})\right),\quad\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}-\bm{w}_{1}^{*}\|_{2}=O\left(\sqrt{s_{1}}(\alpha+\sqrt{\frac{\log p}{n}})\right),$
where $\theta$ is the minimal eigenvalue of $\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}$ and $\alpha=\|\bm{\alpha}\|_{\infty}$.
@@
$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\alpha}=\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}_{\mathcal{A}_{1}}\widehat{\bm{\Sigma}}_{\mathcal{A}_{1},\mathcal{A}_{1}}\bm{w}_{\mathcal{A}_{1}}-\gamma\bm{w}^{\top}_{\mathcal{A}_{1}}\bm{\mu}_{1}+\|\bm{\alpha}_{\mathcal{A}_{1}}\odot\bm{w}^{\top}_{\mathcal{A}_{1}}\|_{1}.$ (2.17)
@@
Under Assumptions (A1)-(A4), we select $\lambda\geq M\sqrt{\frac{s_{1}\log p}{n}}$ for some large constant $M$. Let $\widehat{\bm{w}}^{\text{LLA},\alpha}_{t}$ be the optimum of LLA algorithm initialized by $\widehat{\bm{\bm{w}}}_{t}^{\text{Lasso },\alpha}$. If $s_{1}=O\left(\sqrt{\frac{n}{\log p}}\right)$ and $\|\bm{\alpha}\|_{\infty}\leq C\sqrt{\frac{\log p}{n}}$ for some large constant $C$, $\widehat{\bm{w}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{w}}}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[s_{1},p]$ and some constant $c_{1},c_{2}$. Furthermore,
@@
To ensure the theoretical properties of subsequent proofs, Theorem 3 reasonably assumes $\|\bm{\alpha}\|_{\infty}=O(\sqrt{\log p/n})$, which matches the convergence rate of $\lambda_{\text{Lasso}}$. This differs from the assumption in Section 2.2, where $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A2). This distinction is also evident in our later sections on simulation studies and real data applications, where $\bm{\beta}=0.15\times\mathbf{1}_{p}$ while $\bm{\alpha}=0.001\times\mathbf{1}_{p}$.
@@
Building on the insights from Corollary 1, we derive that under the condition $s_{1}\sqrt{s_{1}\log p}=o(1)$, the estimation errors for in-sample and out-of-sample Sharpe ratios during reallocation periods asymptotically vanish as $n$ approaches infinity. This result highlights the diminishing impact of estimation uncertainty in high-dimensional settings when the sparsity condition is adequately controlled.
@@ 2.3.2 Portfolio Reallocation
$\min_{\bm{w}_{t}}\bm{w}^{\top}_{t}\bm{\Sigma}_{t}\bm{w}_{t}-\gamma\bm{w}^{\top}_{t}\bm{\mu}_{t-1}+\|\bm{\alpha}\odot(\bm{w}_{t}-\bm{w}^{+}_{t-1})\|_{1}\text{ s.t. }\bm{w}_{t}^{\top}\mathbf{1}=1,$
for $t=2,3,...m$, with $\bm{\delta}_{t}=\bm{w}_{t}-\bm{w}^{+}_{t-1}$, which has the equivalent form:
$\min_{\bm{\delta}_{t}}\bm{\delta}^{\top}_{t}\bm{\Sigma}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\bm{\mu}_{t-1}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1},\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0.$ (2.18)
@@
$\bm{\delta}_{t}^{*}=\frac{1}{2}\bm{\Sigma}_{t}^{-1}\left(\gamma\tilde{\bm{\mu}}_{t}-\bm{\alpha}\odot\mathbf{g}+\frac{2+\gamma\mathbf{1}^{\top}\bm{\Sigma}_{t}^{-1}\bm{\alpha}\odot\mathbf{g}-\gamma\mathbf{1}^{\top}\bm{\Sigma}_{t}^{-1}\tilde{\bm{\mu}}_{t}}{\mathbf{1}^{\top}\bm{\Sigma}_{t}^{-1}\mathbf{1}}\mathbf{1}\right)\quad\text{and}\quad\bm{\delta}_{\mathcal{A}^{c}_{t}}^{*}=\mathbf{0},$
where $\tilde{\bm{\mu}}=\bm{\mu}-\frac{2}{\gamma}\bm{\Sigma}_{t}\bm{w}^{+}_{t-1}$ and $\mathbf{g}$ is the vector of sub-derivatives of $\left\|\bm{w}_{t}\right\|_{1}$, i.e., $\mathbf{g}:=\partial\left\|\bm{w}_{t}\right\|_{1}/\partial\bm{w}_{t}$. In practice, the feasible form is
$\min_{\bm{\delta}_{t}}\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\widehat{\bm{w}}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t-1}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}+P_{\lambda}(\bm{\delta}_{t}),\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0.$ (2.19)
@@
$\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}=$ $\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\bm{\delta}^{\top}\mathbf{1}=1}\bm{\delta}^{\top}_{\mathcal{A}_{t}}\widehat{\bm{\Sigma}}_{\mathcal{A}_{t},\mathcal{A}_{t}}\bm{\delta}_{\mathcal{A}_{t}}+2(\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t})_{\mathcal{A}_{t}}\bm{\delta}_{\mathcal{A}_{t}}-\gamma\bm{\delta}^{\top}_{\mathcal{A}_{t}}\bm{\mu}_{\mathcal{A}_{t}}+\|\bm{\alpha}_{\mathcal{A}_{t}}\odot\bm{\delta}_{\mathcal{A}_{t}}\|_{1}.$ (2.20)
@@
Suppose that Assumptions (A1)-(A3) and (A4’) hold and $\|\bm{\alpha}\|_{\infty}=C\sqrt{\frac{\log p}{n}}$ for some large constant $C$. Let $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}$ be the unique optimum of the program $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{\delta}_{t}=0}\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\widehat{\bm{w}}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t-1}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}+\lambda_{Lasso,t}\|\bm{\delta}_{t}\|_{1}$. Then we have $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}-\bm{\delta}_{t}^{*}\|_{1}=O_{p}\left(s_{t}\sqrt{\frac{\log p}{n}}\right)$ and $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}-\bm{\delta}_{t}^{*}\|_{2}=O_{p}\left(\sqrt{s_{t}\frac{\log p}{n}}\right)$. If $s_{t}=O\left(\sqrt{\frac{n}{\log p}}\right)$, we choose $\lambda_{t}\geq M\sqrt{\frac{s_{t}\log p}{n}}$ for some large constant $M$ and let $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ be the optimum of LLA algorithm initialized by $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}.$
@@
    If $\|\bm{w}^{+}_{t-1}\|_{1}<C_{4}$, $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$. Furthermore, $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\alpha}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\log p/n}\right).$
@@
    If $\|\bm{w}^{+}_{t-1,\mathcal{Q}}\|_{1}\leq C_{4}$ and $\bm{w}^{+}_{t-1,\mathcal{Q}^{c}}=\mathbf{0}$ for some set $\mathcal{Q}\subset[1,\ldots,p]$, $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[\max\{s_{t},|\mathcal{Q}|\},p]$ and some constant $c_{1},c_{2}$. Furthermore, $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\alpha}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\log k/n}\right)$.
Remark. Although the conditions and conclusions of Theorem 4 and Theorem 2 appear to be very similar, the differences arise from the distinct assumptions regarding transaction costs. For quadratic transaction costs, we assume only that $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A2). In contrast, for proportional transaction costs, we require $\alpha$ to satisfy a specific convergence rate. This distinction stems from the differing impacts of transaction costs on the optimization problem introduced in Section 2.1.
For quadratic transaction costs, it suffices to verify that the modified covariance matrix continues to satisfy our assumptions, as demonstrated in the Lemma LABEL:Property_of_tilde_bSigma provided in the Supplementary Material. However, for proportional transaction costs, Lemma 2 explicitly illustrates the influence of $\alpha$ on both the convergence rate and consistency. This influence persists in Theorems 3 and 4, where the assumptions regarding transaction costs lead to similar conclusions despite the differences in the underlying hypotheses.
Analogous to Corollary 2, it can be established that if the sparsity level and dimensionality satisfy $k_{t}\sqrt{k_{t}\log p}=o(1)$, the estimation errors for both in-sample and out-of-sample Sharpe ratios during reallocation periods converge to zero as $n\to\infty$.

### v2

@@ 2.4.1 First Portfolio Construction
$\min_{\bm{w}_{1}}\ \bm{w}^{\top}_{1}\bm{\Sigma}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\bm{\mu}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1},\text{ s.t. }\bm{w}_{1}^{\top}\mathbf{1}=1,\ \|\bm{w}_{1}\|_{0}\leq s_{0},$
where $\bm{\alpha}=(\alpha_{1},...,\alpha_{p})^{\top},\alpha_{j}>0$ is the cost parameter, which is similar to $\bm{\beta}$ in Section 2.3. By Lagrange multiplier method and the definition of $\bm{w}^{*}_{1}$ and $\mathcal{A}_{1}$, solving for $\bm{w}_{1}^{*}$ yields
$\bm{w}_{\mathcal{A}_{1}}^{*}=\frac{1}{2}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\left(\gamma\bm{\mu}_{\mathcal{A}_{1}}-\bm{\alpha}\odot\mathbf{g}_{\mathcal{A}_{1}}+\frac{2-\gamma\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\bm{\mu}_{\mathcal{A}_{1}}+\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}(\bm{\alpha}\odot\mathbf{g}_{\mathcal{A}_{1}})}{\mathbf{1}^{\top}\bm{\Sigma}^{-1}_{\mathcal{A}_{1},\mathcal{A}_{1}}\mathbf{1}}\mathbf{1}\right)\quad\text{and}\quad\bm{w}^{*}_{\mathcal{A}^{c}_{1}}=\mathbf{0}.$
where $\mathbf{g}$ is the vector of sub-derivatives of $\left\|\bm{w}_{1}\right\|_{1}$, i.e., $\mathbf{g}:=\partial\left\|\bm{w}_{1}\right\|_{1}/\partial\bm{w}_{1}$, consisting of elements which are 1 or 1-1 in case $w_{j}>0$ or $w_{j}<0$, respectively, or $\in[-1,1]$ in case $w_{j}=0$. Similar to Lemma 1 in Section 2.3, Lemma 3 shows the consistency of the CAPE-L $\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}$ under proportional transaction cost.
@@
$\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{w}_{1}=1}\bm{w}_{1}^{\top}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}_{1}^{\top}\widehat{\bm{\mu}}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}+\lambda_{\text{Lasso},1}\|\bm{w}_{1}\|_{1},$ (2.17)
with $\lambda_{\text{Lasso},1}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$, then with probability at least $1-c_{1}p^{-c_{2}}$ for some constants $c_{1},c_{2}$, we have
$\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}-\bm{w}_{1}^{*}\|_{1}=O\left(s_{1}\|\bm{\alpha}\|_{\infty}+s_{1}\sqrt{\frac{\log p}{n}}\right)\quad\text{and}\quad\|\widehat{\bm{\bm{w}}}_{1}^{\text{Lasso},\alpha}-\bm{w}_{1}^{*}\|_{2}=O\left(\sqrt{s_{1}}\|\bm{\alpha}\|_{\infty}+\sqrt{\frac{s_{1}\log p}{n}}\right).$
@@
$\widehat{\bm{\bm{w}}}_{1}^{\text{oracle},\alpha}=\mathop{\arg\min}_{\bm{\bm{w}}_{\mathcal{A}_{1}^{c}}=\mathbf{0},\bm{w}^{\top}\mathbf{1}=1}\bm{w}^{\top}_{1}\widehat{\bm{\Sigma}}_{1}\bm{w}_{1}-\gamma\bm{w}^{\top}_{1}\widehat{\bm{\mu}}_{1}+\|\bm{\alpha}\odot\bm{w}_{1}\|_{1}.$ (2.18)
@@
Under Assumptions (A1)-(A3) and $\|\bm{\alpha}\|_{\infty}\leq C\sqrt{\frac{\log p}{n}}$ for some large constant $C$, we select $\lambda\geq M\sqrt{\frac{s_{1}\log p}{n}}$ for some large constant $M$. Let $\widehat{\bm{w}}^{\text{LLA},\alpha}_{t}$ be the optimum of LLA algorithm 1 with SCAD penalty, initialized at $\widehat{\bm{\bm{w}}}_{t}^{\text{Lasso },\alpha}$. If $s_{1}=O\left(\sqrt{\frac{n}{\log p}}\right)$ and $\left\|\bm{w}^{*}_{1}\right\|_{\min}>(a+1)\lambda_{1}$, $\widehat{\bm{w}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{w}}}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for $\forall k\in[s_{1},p]$ and some constant $c_{1},c_{2}$. Furthermore,
@@
To ensure the theoretical properties of subsequent proofs, Theorem 3 reasonably assumes $\|\bm{\alpha}\|_{\infty}=O(\sqrt{\log p/n})$, which matches the convergence rate of $\lambda_{\text{Lasso}}$. This differs from the assumption in Section 2.3, where $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A1).
@@
As in Corollary 1, we can show that under the condition $s_{1}\sqrt{s_{1}\log p}=o(1)$, the estimation errors for in-sample and out-of-sample Sharpe ratios during reallocation periods asymptotically vanish as $n$ approaches infinity. Since the conclusion is identical to that of Corollary 1, we omit the formal statement here. This result highlights the diminishing impact of estimation uncertainty in high-dimensional settings when the sparsity condition is adequately controlled.
@@ 2.4.2 Portfolio Reallocation
$\min_{\bm{\delta}_{t}}\ \bm{\delta}^{\top}_{t}\bm{\Sigma}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\bm{\Sigma}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\bm{\mu}_{t}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1},\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0,\ \|\bm{\delta}_{t}\|_{0}\leq s_{0}.$
@@
$\bm{\delta}_{\mathcal{A}_{t}}^{*}=\frac{1}{2}\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\left(\gamma\bm{\mu}_{\mathcal{A}_{t}}-2(\bm{\Sigma}_{t}\bm{w}^{+}_{t-1})_{\mathcal{A}_{t}}-\bm{\alpha}\odot\mathbf{g}_{\mathcal{A}_{t}}+\frac{2+\gamma\mathbf{1}^{\top}\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\bm{\alpha}\odot\mathbf{g}_{\mathcal{A}_{t}}-\gamma\mathbf{1}^{\top}\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\bm{\mu}_{\mathcal{A}_{t}}}{\mathbf{1}^{\top}\bm{\Sigma}_{\mathcal{A}_{t},\mathcal{A}_{t}}^{-1}\mathbf{1}}\mathbf{1}\right)\ \text{and}\ \bm{\delta}_{\mathcal{A}^{c}_{t}}^{*}=\mathbf{0},$
where $\mathbf{g}$ is the vector of sub-derivatives of $\left\|\bm{w}_{t}\right\|_{1}$, i.e., $\mathbf{g}:=\partial\left\|\bm{w}_{t}\right\|_{1}/\partial\bm{w}_{t}$. In practice, the feasible form is
$\min_{\bm{\delta}_{t}}\ \bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}+P_{\lambda}(\bm{\delta}_{t}),\text{ s.t. }\bm{\delta}_{t}^{\top}\mathbf{1}=0.$
@@
$\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}=$ $\mathop{\arg\min}_{\bm{\bm{\delta}}_{\mathcal{A}_{t}^{c}}=\mathbf{0},\bm{\delta}_{t}^{\top}\mathbf{1}=0}\bm{\delta}^{\top}_{t}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}_{t-1}^{+\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}^{\top}_{t}\widehat{\bm{\mu}}_{t}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}.$ (2.19)
Lemma 4.
Suppose that Assumptions (A1)–(A3) hold, and that $\|\bm{\alpha}\|_{\infty}=C\sqrt{\frac{\log p}{n}}$ for some large constant $C$. Let $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}$ denote the unique solution to the following optimization problem:
$\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}=\mathop{\arg\min}_{\mathbf{1}^{\top}\bm{\delta}_{t}=0}\bm{\delta}_{t}^{\top}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}+2\bm{w}^{+\top}_{t-1}\widehat{\bm{\Sigma}}_{t}\bm{\delta}_{t}-\gamma\bm{\delta}_{t}^{\top}\widehat{\bm{\mu}}_{t-1}+\|\bm{\alpha}\odot\bm{\delta}_{t}\|_{1}+\lambda_{\text{Lasso},t}\|\bm{\delta}_{t}\|_{1}.$ (2.20)
with $\lambda_{\text{Lasso},t}=M\sqrt{\frac{\log p}{n}}$ for some large constant $M$. Then with probability at least $1-c_{1}p^{-c_{2}}$ for some constant $c_{1},c_{2}$, we have
$\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}-\bm{\delta}_{t}^{*}\|_{1}=O_{p}\left(s_{t}\sqrt{\frac{\log p}{n}}\right)\quad\text{and}\quad\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}-\bm{\delta}_{t}^{*}\|_{2}=O_{p}\left(\sqrt{\frac{s_{t}\log p}{n}}\right).$
@@
Suppose that Assumptions (A1)–(A3) hold and that $\|\bm{\alpha}\|_{\infty}=C\sqrt{\frac{\log p}{n}}$ for some large constant $C$. If $s_{t}=O\left(\sqrt{\frac{n}{\log p}}\right)$ and $\|\bm{\delta}_{t}^{*}\|_{\min}>(a+1)\lambda_{t}$, choose $\lambda_{t}\geq M\sqrt{\frac{s_{t}\log p}{n}}$ for some large constant $M$. Let $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ be the optimum of the LLA algorithm 2 with SCAD penalty, initialized at $\widehat{\bm{\bm{\delta}}}_{t}^{\text{Lasso},\alpha}$ as defined in (2.20). Then, the following holds:
@@
    If $\|\bm{w}^{+}_{t-1}\|_{1}<C_{0}$, then $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}p^{-c_{2}}$ for some constants $c_{1},c_{2}>0$. Furthermore,
    $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\alpha}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log p}{n}}\right).$
@@
    If $\|\bm{w}^{+}_{t-1,\mathcal{Q}}\|_{1}\leq C_{0}$ and $\bm{w}^{+}_{t-1,\mathcal{Q}^{c}}=\mathbf{0}$ for some set $\mathcal{Q}\subset[1,\ldots,p]$ and absolute constant $C_{0}$, then $\widehat{\bm{\delta}}^{\text{LLA},\alpha}_{t}$ converges to $\widehat{\bm{\bm{\delta}}}_{t}^{\text{oracle},\alpha}$ after two iterations with probability at least $1-c_{1}k^{-c_{2}}$ for all $k\in[\max\{s_{t},|\mathcal{Q}|\},p]$ and some constants $c_{1},c_{2}>0$. Furthermore,
    $\|\widehat{\bm{\bm{\delta}}}_{t}^{\text{LLA},\alpha}-\bm{\delta}_{t}^{*}\|_{\infty}=O\left(\sqrt{\frac{\log k}{n}}\right).$
Remark 2.
Although the conditions and conclusions of Theorem 4 and Theorem 2 appear to be very similar, the differences arise from the distinct assumptions regarding transaction costs. For quadratic transaction costs, we assume only that $\|\bm{\beta}\|_{\infty}\ll 1/C_{2}$ with $C_{2}$ defined in Assumption (A1). In contrast, for proportional transaction costs, we require $\alpha$ to satisfy a specific convergence rate. This distinction stems from the differing impacts of transaction costs on the optimization problem introduced in Section 2.2.
For quadratic transaction costs, it suffices to verify that the modified covariance matrix continues to satisfy our assumptions, as demonstrated in the Supplementary Material. However, for proportional transaction costs, Lemma 3 explicitly illustrates the influence of $\alpha$ on both the convergence rate and consistency. This influence persists in Theorems 3 and 4, where the assumptions regarding transaction costs lead to similar conclusions despite the differences in the underlying hypotheses.
Analogous to Corollary 2, it can be established that if the sparsity level and dimensionality satisfy $k_{t}\sqrt{k_{t}\log p}=o(1)$, the estimation errors for both in-sample and out-of-sample Sharpe ratios during reallocation periods converge to zero as $n\to\infty$. Since the conclusion is identical to that of Corollary 2, we omit the formal statement here.